OpenStudy (anonymous):

What is the exact value of cos 225° cos 75°?

OpenStudy (anonymous):

cos 225 = -sqrt2/2

OpenStudy (anonymous):

cos 75=cos(30 + 45) =cos30 cos45 - sin30 sin45 =sqrt3 - 1/2 sqrt2